# Tight contact structures without symplectic fillings are everywhere

### 27 Feb 2024 v4

@@ 2.4. Capping Giroux domains
In [BGM22, Theorem B] certain homological obstructions to the symplectic fillability of Bourgeois contact manifolds were obtained. The argument of the proof uses a capping construction of Massot-Niederkrüger-Wendl [MNW13]. More precisely, whenever one finds an embedded Giroux domain of the form $G=X\times\mathbb{S}^{1}$, one can perform a blow down construction as in [MNW13], which topologically consists in removing the interior of $G$ and collapsing all circle fibers at the remaining boundary to points. By [MNW13, Section 5.1] the resulting smooth manifold $M_{bd}$ (where $bd$ stands for blown down) carries a contact structure that is well-defined (up to isotopy). The following special case of [MNW13, Theorem 6.1] then says that this is achievable through a cobordism:
@@
Topologically, this cobordism is obtained by attaching $X\times\mathbb{D}^{2}$ on top of $X\times\mathbb{S}^{1}$ (and smoothing corners), where the latter is seen as the positive boundary of $[0,1]\times X\times\mathbb{S}^{1}$. We emphasize that this cobordism is not Liouville, as there are only local Liouville vector fields near the boundary in general.
@@
Let $M=V\times\mathbb{S}^{1}$ be a contact manifold be a contact manifold with an $\mathbb{S}^{1}$-invariant contact structure and convex decomposition
@@
Then there is a strong symplectic cobordism from $M$ to the contact boundary of $V_{+}\times\mathbb{D}^{2}$. Furthermore, if $c_{1}(M)$ is torsion, then the same is true for the cobordism.
@@ 2.5. Allowing surgeries on a Giroux domain
We now wish to state a generalisation of Corollary 2.12, where we allow contact surgeries in one part of the convex decomposition.
@@
Consider an $\mathbb{S}^{1}$-equivariant contact structure on $M=V^{2n}\times\mathbb{S}^{1}$, with induced convex splitting $V=V_{+}\cup\overline{V}_{-}$. Suppose now that $(M^{\prime},\xi^{\prime})$ is obtained from the former via a sequence of subcritical/flexible surgeries on isotropic/loose Legendrian spheres contained in the complement of $V_{-}\times\mathbb{S}^{1}$.
@@
Moreover, if $V_{+}$ is Weinstein, there is a (smooth) copy of $V_{+}$ inside $M_{+}^{\prime}$, such that the following holds:
  • $V_{+}\subset M_{+}^{\prime}$ can be isotoped to lie in the negative end $M^{\prime}$ of the cobordism $W_{cap}$;
[…]
@@
For the second part of the statement, first notice that, because of the dimensions involved, the attaching regions of the surgery handles that give $M_{+}^{\prime}$ starting from $OB(V_{+},\operatorname{Id})$ can be made smoothly disjoint from the $n$-skeleton $V_{+}^{(n)}$ of a page $V_{+}$ in the open book by general position. In particular, by the Weinstein assumption on $V_{+}$, one naturally has a copy of $N(V_{+}^{(n)})\simeq V_{+}$ inside $M_{+}^{\prime}$. What is more, as $M_{+}^{\prime}$ is topologically obtained from $M^{\prime}$ by attaching a handle $V_{-}\times\mathbb{D}^{2}$, i.e. Morse handles of indices between $1$ and $n+1$, the subset $V_{+}^{(n)}\subset M_{+}^{\prime}$ can be isotoped down to the negative boundary $M^{\prime}$ of $W_{cap}$, and the same is true for $N(V_{+}^{(n)})\simeq V_{+}\subset M_{+}^{\prime}$. The final claim follows from the fact that it is obtained by performing a certain number, say $k$, of contact surgeries on $OB(V_{+},\operatorname{Id})$, which can at most kill $k$ generators in the homology group of the corresponding degree. ∎
@@
[…]
@@ 3. An algebraic perspective on tightness
By the work of Bourgeois and van-Koert [BvK10], the contact homology, defined rigorously on the homology level by Pardon [Par19] and Bao-Honda [BH23], vanishes for overtwisted manifolds. Motivated from this, Bourgeois and Niederkrüger [BN10] introduced the notion of algebraically overtwisted manifolds as follows.
@@
    Let $W$ be a Liouville cobordism with convex boundary $\partial_{+}W$ and concave boundary $\partial_{-}W$. If $\partial_{+}W$ is algebraically overtwisted then so is $\partial_{-}W$.
@@
    The following contact manifolds are algebraically tight.
@@
      Contact manifolds with vanishing rational first Chern class and a contact form $\alpha$, such that there is no contractible Reeb orbit of SFT degree $1$;
@@
The first claim follows from [BvK10] and the main result of [CMP19]. Although [BvK10] uses $\mathbb{Q}$-coefficients, the equivalence with using $\Lambda$-coefficients can be found in [MZ20, Proposition 3.12]. The second claim follows from the fact that $\operatorname{CH}(Y_{1}\sqcup Y_{2})=\operatorname{CH}(Y_{1})\otimes\operatorname{CH}(Y_{2})$. The third claim follows from the functoriality of contact homology, i.e. we have a unital algebra map $\operatorname{CH}(\partial_{+}W)\to\operatorname{CH}(\partial_{-}W)$. The fourth claim follows from the fact that $\operatorname{CH}(Y;\Lambda)=0$ is an obstruction to the existence of augmentations, hence existence of strong fillings, see e.g. [MZ20, Proposition 2.11, 3.13]. Conditions in (a) and (b) of the fifth claim mean that $1\neq 0\in\operatorname{CH}(Y)$, which implies algebraic tightness by homotopy class or degree reasons. Roughly speaking, (c) is a special case of (b). (Recall that $k-$ADC contact structures have torsion $c_{1}$ by definition.) However, to cope with the asymptotic property in the definition of $1$-ADC, we can argue as follows. If $\operatorname{CH}(Y)=0$, then for any fixed contact form $\alpha$, there exists a positive number $A$, such that $1=0\in\operatorname{CH}^{<A}(Y,\alpha)$, i.e. the homology of the sub-complex generated by $\alpha$-orbits with contact action at most $A$. By the $1$-ADC condition, we can find a contact form $\alpha_{0}<\alpha$, such that all $\alpha_{0}$-Reeb orbits with action smaller than $A$ have SFT degree at least $2$. Then the functoriality $\operatorname{CH}^{<A}(Y,\alpha)\to\operatorname{CH}^{<A}(Y,\alpha_{0})$ implies that $1=0\in\operatorname{CH}^{<A}(Y,\alpha_{0})$, which is impossible due to degree reasons. ∎
@@
Starting with an open book whose binding is $1$-ADC and simply connected, according to Lemma 2.8 the associated Bourgeois contact structure is also $1$-ADC (in fact $3$-ADC), hence algebraically tight. As special cases, Bourgeois manifolds in Example 2.10 are algebraically tight. Then after contact surgeries, the contact manifold remains algebraically tight, hence tight.
@@
One of the most fundamental questions in symplectic and contact topology is whether the boundary between rigidity and flexibility is captured by pseudo-holomorphic curves. In this framework, it was shown by Avdek [Avd20] that algebraic overtwistedness in dimension $3$ does not imply overtwistedness. Avdek used another tightness criterion from Heegaard-Floer homology, which is still based on pseudo-holomorphic curves. Unlike the situation in dimension $3$, in higher dimensions, it seems that algebraic tightness is the only currently available criterion to ensure tightness.
@@
$5$-dimensional Bourgeois contact structures $BO(\Sigma,\psi)$ are algebraically tight if $\Sigma$ is non-sporadic (i.e. not a sphere with $3$ or less punctures).
@@
The hypothesis of page non-sporadic gives that there is a strong symplectic cobordisms whose negative end consists of hypertight contact manifolds [BGM22, Theorem 3.1 and Corollary 4.1]. Now the symplectic form on this cobordism is exact (hence no sphere bubbling can occur), however the given global primitive $1$-form is not compatible with the contact structure on the negative end (this cobordism is called pseudo-Liouville in [BGM22]). In terms of functoriality in SFT, a priori this leads to deformations of the algebraic structure on the negative boundary by counting holomorphic caps in the cobordism. However, symplectic caps for this specific cobordism can be ruled out as shown in the proof of [BGM22, Lemma 5.1]. Hence the usual functoriality for contact homology holds (over Novikov coefficients), and the positive boundary is also algebraically tight, as hypertight contact manifolds are algebraically tight. ∎
@@
4. Obstructions to strong fillings
4.1. Homological restrictions on symplectic fillings
@@
Theorem 4.1.
Let $W_{0}$ be a $(2n+2)$-dimensional subcritical Weinstein domain with $c_{1}(W_{0})=0$, and $M$ any contact manifold. If $W$ is a strong filling of $M\#\partial W_{0}$, then the kernel of $H_{n}(\partial W_{0};\mathbb{Q})\to H_{n}(W_{0};\mathbb{Q})$ contains the kernel of $H_{n}(\partial W_{0};\mathbb{Q})\stackrel{{\scriptstyle 0\oplus\operatorname{Id}}}{{\hookrightarrow}}H_{n}(M\#\partial W_{0};\mathbb{Q})\to H_{n}(W;\mathbb{Q})$.
Moreover, the same conclusion holds under the hypothesis that $W_{0}$ is a flexible Weinstein domain of dimension $6$, provided that the contact structure on $M$ has vanishing $c_{1}$.
@@
Theorem 4.1 is stated in this way only for the purpose of constructing exotic contact structures on spheres in this paper, but in fact a stronger version of this result, giving homological information on the fillings in some other degrees for $W_{0}$ flexible, can also be obtained by similar arguments. We also point out that, when $M=\emptyset$ and $W_{0}$ is subcritical, one can use a strategy of filling by pseudo-holomorphic curves similar to [BGZ19] to obtain a similar result, which was used by the first three authors [BGM22, Theorem B] to obstruct strong fillings.
Theorem 4.1 should be understood in the context of the classification of fillings of special families of contact manifolds in higher dimensions. Taking the celebrated Eliashberg–Floer–McDuff theorem [McD91] as a starting point, many contact manifolds are expected to have unique Liouville/symplectically aspherical fillings, at least up to diffeomorphism. Up until now, results in this direction mainly apply to the contact boundary of a split manifold of the form $V\times\mathbb{D}^{2}$, up to subcritical/flexible surgeries; we refer to [OV12, BGZ19, Zho21b, Zho22a, BGM22, GKZ23].
If one considers the more general class of strong fillings, these are no longer unique, as one can take a blow-up of any given filling without affecting the boundary. However, it is plausible that all the ambiguity only comes from birational surgeries like blow-ups, which will increase the complexity of the topology of the filling; see [Zho21b, §8] for supporting evidence. Moreover, Eliashberg [Eli90] showed in dimension $3$, that if $W$ is a symplectic filling of $Y\#Y^{\prime}$, then $W$ is obtained from attaching a $1$-handle to a symplectic filling $W^{\prime}$ of $Y_{1}\sqcup Y_{2}$ (which might be connected, cf. [Bow12]). Although the higher dimensional analogue has not been established, partial generalizations were obtained by Ghiggini-Niederkrüger-Wendl [GNW16]. Incarnations of these expectations at the level of homology lead to our obstructions to strong fillings.

### 14 Mar 2026 v5

@@ 2.4. Capping Giroux domains
In [BGM22, Theorem B], certain homological obstructions to the symplectic fillability of Bourgeois contact manifolds were obtained. The argument of the proof uses a capping construction of Massot-Niederkrüger-Wendl [MNW13]. More precisely, whenever one finds an embedded Giroux domain of the form $G=X\times\mathbb{S}^{1}$, one can perform a blow down construction as in [MNW13], which topologically consists of removing the interior of $G$ and collapsing all circle fibers at the remaining boundary to points. By [MNW13, Section 5.1], the resulting smooth manifold $M_{bd}$ (where $bd$ stands for blown down) carries a contact structure that is well-defined (up to isotopy). The following special case of [MNW13, Theorem 6.1] then says that this is achievable through a cobordism:
@@
Topologically, this cobordism is obtained by attaching $X\times\mathbb{D}^{2}$ on top of $X\times\mathbb{S}^{1}$ (and smoothing corners), where the latter is seen as the positive boundary of $0,1]\times X\times\mathbb{S}^{1}$. We emphasize that this cobordism is not Liouville, as there are only local Liouville vector fields near the boundary in general.
@@
Let $M=V\times\mathbb{S}^{1}$ be a contact manifold with an $\mathbb{S}^{1}$-invariant contact structure and convex decomposition
@@
Then there is a strong symplectic cobordism from $M$ to the contact boundary of $V_{+}\times\mathbb{D}^{2}$. Furthermore, if $c_{1}(\xi)$ is torsion and $V_{-}$ is connected, then the same is true for the first Chern class of the cobordism.
The last sentence in the above statement follows from the following observation. The cobordism $W$ above topologically corresponds to attaching the handle $V_{-}\times\mathbb{D}^{2}$ to $M\times[0,1]$ along $\overline{V}_{-}\times\mathbb{S}^{1}\subset M\times\{1\}$. By the long exact sequence in homology associated to the pair $(W,M)$, we have an exact sequence
$H_{2}(M;\mathbb{Q})\to H_{2}(W;\mathbb{Q})\to H_{2}(W,M;\mathbb{Q})\to H_{1}(M;\mathbb{Q}).$
Since $H_{2}(W,M;\mathbb{Q})\simeq H_{2}(V_{-}\times\mathbb{D}^{2},V_{-}\times\mathbb{S}^{1};\mathbb{Q})\simeq H_{2}(\{p\}\times\mathbb{D}^{2},\{p\}\times\mathbb{S}^{1};\mathbb{Q})\simeq H_{1}(\{p\}\times\mathbb{S}^{1};\mathbb{Q})$ when $V_{-}$ is connected, which injects into $H_{1}(M=V\times\mathbb{S}^{1};\mathbb{Q})$ for $p\in V_{-}$. Therefore, the inclusion of the negative end $M\hookrightarrow W$ induces a surjection on $H_{2}(\cdot;\mathbb{Q})$. In particular, $c_{1}(W)\in H^{2}(W;\mathbb{Q})$ is determined by its restriction to the negative boundary of $W$, which just coincides with $c_{1}(\xi)$. Hence, the latter being torsion implies that the former is torsion as well.
@@ 2.5. Allowing surgeries on a Giroux domain
We now wish to state a generalization of Corollary 2.12, where we allow contact surgeries in one part of the convex decomposition.
@@
Consider an $\mathbb{S}^{1}$-equivariant contact structure on $M=V^{2n}\times\mathbb{S}^{1}$, with induced convex splitting $V=V_{+}\cup\overline{V}_{-}$. Suppose now that $(M^{\prime},\xi^{\prime})$ is obtained from the former via a sequence of contact surgeries on isotropic (possibly Legendrian) spheres contained in the complement of $V_{-}\times\mathbb{S}^{1}$.
@@
Moreover, if $V_{+}$ is Weinstein, there is a (smooth) copy $V_{+}^{\prime}$ of $V_{+}$ inside $M_{+}^{\prime}$, such that the following holds:
  • $V_{+}^{\prime}\subset M_{+}^{\prime}$ can be isotoped to lie in the negative end $M^{\prime}$ of the cobordism $W_{cap}$;
  • the image of the map induced by the natural inclusion $H_{q}(V^{\prime}_{+})\to H_{q}(M^{\prime}_{+})$, has dimension at least $\dim H_{q}(V^{\prime}_{+})-\#(q\text{-surgeries})$.
@@
For the second part of the statement, we first point out that the dimension of the ambient manifold is $2n+1$, while $V_{+}$ is homotopy equivalent to its $n$-skeleton, and the dimension of the attaching spheres for the surgeries is at most $n$. In particular, up to perturbing the attaching spheres of the surgery handles that give $M_{+}^{\prime}$ starting from $OB(V_{+},\operatorname{Id})$ among isotropic/Legendrian submanifolds, we can assume that they are disjoint from the $n$-skeleton $V_{+}^{(n)}$ of a page $V_{+}$ in the open book by general position. In particular, one naturally has a copy of a neighborhood of $V_{+}^{(n)}$, which is homotopy equivalent to $V_{+}$, inside $M_{+}^{\prime}$; we denote it $V_{+}^{\prime}\simeq V_{+}$.
What is more, as $M_{+}^{\prime}$ is topologically obtained from $M^{\prime}$ by attaching a handle $V_{-}\times\mathbb{D}^{2}$, and the subset $V_{+}^{\prime}\subset M_{+}^{\prime}$ is disjoint from it as well as from the surgery region, then $V_{+}^{\prime}$ lives in a region where the cobordism $W_{cap}$ is a trivial product and therefore $V_{+}^{\prime}$ can be isotoped down to the negative boundary $M^{\prime}$ of $W_{cap}$.
The final claim follows from the fact that $M_{+}^{\prime}$ is obtained by performing a certain number, say $k$, of contact surgeries on $OB(V_{+},\operatorname{Id})$, which can at most kill $k$ generators in the homology group of the corresponding degree. ∎
@@
Figure 1. The blow-down handle $V_{-}\times D^{2}$ in green, attached on top of $\{1\}\times V_{-}\times\mathbb{S}^{1}\subset\{1\}\times M$ seen as the boundary of the trivial cobordism $0,1]\times M$. The purple region depicts the handle attachment that gives $M^{\prime}$ from $M$, attached on the region $V_{+}\times\mathbb{S}^{1}\subset M$. In red the contact manifold $M^{\prime}_{+}$ resulting from both blow-down and contact surgeries.
@@ 3. An algebraic perspective on tightness
Bourgeois and van-Koert [BvK10] showed that contact homology, which has now been defined rigorously on the homology level by Pardon [Par19] and Bao-Honda [BH23], vanishes for overtwisted contact manifolds. Motivated by this, Bourgeois and Niederkrüger [BN10] introduced the notion of algebraically overtwisted contact manifolds as follows.
@@
    Let $W$ be a Liouville cobordism with convex boundary $\partial_{+}W$ and concave boundary $\partial_{-}W$. If $\partial_{+}W$ is algebraically overtwisted, then so is $\partial_{-}W$.
@@
    The following contact manifolds are algebraically tight:
@@
      contact manifolds with vanishing rational first Chern class and a contact form $\alpha$ such that there is no contractible Reeb orbit of SFT degree $1$;
@@
The first claim follows from [BvK10] and the main result of [CMP19]. Although [BvK10] uses $\mathbb{Q}$-coefficients, the claim is also true with $\Lambda$-coefficients (i.e. vanishing of contact homology is independent of these choices), as the power of the Novikov variable in the Novikov field is determined by the periods of Reeb orbits (see e.g. [MZ25, Proof of Corollary 3.15]). The second claim follows from the fact that $\operatorname{CH}(Y_{1}\sqcup Y_{2})=\operatorname{CH}(Y_{1})\otimes\operatorname{CH}(Y_{2})$. The third claim follows from the functoriality of contact homology, i.e. we have an unital algebra map $\operatorname{CH}(\partial_{+}W)\to\operatorname{CH}(\partial_{-}W)$. The fourth claim follows from the fact that $\operatorname{CH}(Y;\Lambda)=0$ is an obstruction to the existence of augmentations, hence the existence of strong fillings, see e.g. [MZ25, Corollary 3.15]. Conditions in (a) and (b) of the fifth claim imply that $1\neq 0\in\operatorname{CH}(Y)$ by homotopy class or degree reasons, hence $\operatorname{CH}(Y)\neq 0$. Roughly speaking, (c) is a special case of (b). (Recall that $k$-ADC contact structures have torsion $c_{1}$ by definition.) However, to cope with the asymptotic property in the definition of $1$-ADC, we can argue as follows. If $\operatorname{CH}(Y)=0$, then for any fixed contact form $\alpha$, there exists a positive number $A$, such that $1=0\in\operatorname{CH}^{<A}(Y,\alpha)$, i.e. the homology of the sub-complex generated by $\alpha$-orbits with contact action at most $A$. By the $1$-ADC condition, we can find a contact form $\alpha_{0}<\alpha$, such that all $\alpha_{0}$-Reeb orbits with action smaller than $A$ have SFT degree at least $2$. Then the functoriality $\operatorname{CH}^{<A}(Y,\alpha)\to\operatorname{CH}^{<A}(Y,\alpha_{0})$ implies that $1=0\in\operatorname{CH}^{<A}(Y,\alpha_{0})$, which is impossible due to degree reasons. ∎
@@
Starting with an open book whose binding is (1)(-1)-ADC and simply connected, according to Lemma 2.8, the associated Bourgeois contact structure is $1$-ADC, hence algebraically tight. As special cases, Bourgeois manifolds in Example 2.10 are algebraically tight. Then, after contact surgeries, the contact manifold remains algebraically tight, hence tight.
@@
One of the most fundamental questions in symplectic and contact topology is whether the boundary between rigidity and flexibility is captured by pseudo-holomorphic curves. In this framework, it was shown by Avdek [Avd23] that algebraic overtwistedness in dimension $3$ does not imply overtwistedness. Avdek used another tightness criterion from Heegaard-Floer homology, which is still based on pseudo-holomorphic curves. Unlike the situation in dimension $3$, in higher dimensions, it seems that algebraic tightness is the only currently available criterion to ensure tightness.
@@
55-dimensional Bourgeois contact structures $BO(\Sigma,\psi)$ are algebraically tight if $\Sigma$ is non-sporadic (i.e. not a sphere with $3$ or less punctures).
@@
The hypothesis of non-sporadic page gives that there is a strong symplectic cobordism whose negative end consists of hypertight contact manifolds [BGM22, Theorem 3.1 and Corollary 4.1]. Now the symplectic form on this cobordism is exact (hence no sphere bubbling can occur), however the given global primitive $1$-form is not compatible with the contact structure on the negative end (this cobordism is called pseudo-Liouville in [BGM22]). In terms of functoriality in SFT, a priori this leads to deformations of the algebraic structure on the negative boundary by counting holomorphic caps in the cobordism, see [MZ23]. However, holomorphic caps for this specific cobordism can be ruled out as shown in the proof of [BGM22, Lemma 5.1]. Hence the usual functoriality for contact homology holds (over Novikov coefficients), and the positive boundary is also algebraically tight, as hypertight contact manifolds are algebraically tight. ∎
@@
4. Homological restrictions on symplectic fillings
We begin by giving some context to our main technical result that puts restrictions on the homology of strong symplectic fillings of certain contact manifolds. For the moment we will avoid technicalities and refer to the following sections for more details.
Uniqueness results
Eliashberg–Floer–McDuff theorem [McD91] showed that any Liouville filling of the standard sphere of dimension at least $5$ is diffeomorphic to a ball. In order to achieve this, one considers certain moduli spaces of holomorphic curves that serve as tools to determine the topology of the filling. In particular, one shows that the homology of the filling is fixed by that on the boundary. Uniqueness results of this kind have since been obtained for more general classes of manifolds and apply to fillings of the contact boundary of a split manifold of the form $V\times\mathbb{D}^{2}$, or to flexible Weinstein domains; cf. [OV12, BGZ19, Zho21b, Zho23b, BGM22, GKZ23].
If one considers the more general class of strong fillings, these are no longer unique, as one can take a blow-up of any given filling without affecting the boundary. In particular, one expects a lower bound on the homological complexity of a strong filling coming from a standard Liouville filling.
Symplectic cohomology and fillings
In contrast to arguments based on the Eliashberg-Floer-McDuff result, there is another approach that uses symplectic cohomology. It is well known that the symplectic cohomology of any subcritical Weinstein domain vanishes [Cie02a], and this was extended to flexible Weinstein domains with the advent of Murphy’s notion of loose Legendrians [BEE12, MS18]. In particular, for a flexible filling $W_{0}$, the tautological sequence yields
$\ldots\to SH^{*}(W_{0})=0\to SH_{+}^{*}(W_{0})\stackrel{{\scriptstyle\cong}}{{\to}}H^{*}(W_{0})[1]\to SH^{*+1}(W_{0})=0\to\ldots$
so that the positive symplectic cohomology is isomorphic to the ordinary cohomology of the filling with a grading shift.
Dynamics gives contact invariants and homological constraints
A priori, positive symplectic cohomology might depend on the choice of the Liouville filling. Bourgeois and Oancea [BO09a] showed that there is a way to understand the role of fillings in positive symplectic cohomology in terms of augmentations/counting rigid holomorphic planes from the filling. As a result, Cieliebak and Oancea [CO18] utilised a neck-stretching argument to show that positive symplectic cohomology does not depend on the choice of filling $W$ (with mild topological assumptions) given the contact boundary is index positive as in Definition 2.5. The same result was generalized to the boundary of flexible Weinstein domains with vanishing first Chern class by Lazarev [Laz20a], as the contact boundary is ADC (cf. Definition 2.4). More precisely, under the index-positive/ADC assumption, the Floer differential in positive symplectic cohomology for suitable choices of contact forms and almost complex structures “lives” entirely in the end of the filling. In this way, one gets that the homology of a filling of a flexibly fillable contact manifold (of vanishing first Chern class) is an invariant.
In fact, much more is true: one can consider the restriction of the connecting homomorphism to the boundary, which also does not depend on the filling for ADC manifolds.
$SH_{+}^{*}(W)\stackrel{{\scriptstyle\delta}}{{\to}}H^{*}(W)[1]\to H^{*}(\partial W)[1].$
This then gives an important perspective on this invariant, namely we can understand the image of $H^{*}(W)\to H^{*}(\partial W)$ of any Liouville filling via the invariant map $\delta_{\partial}:SH_{*}^{+}(W)\to H^{*}(\partial W)[1]$ in a standard filling. This viewpoint was used to deduce topological information of Liouville fillings in [Zho21b]. Moreover, in the split case $W=V\times\mathbb{D}^{2}$, similar results hold without any assumptions on the first Chern class [Zho23b].
General strong fillings
In order to extend these ideas to strong fillings and to provide filling obstructions to obtain Theorem A, one needs to extend the above discussion to this more general class of fillings and localize the properties under connected sums. The former concern requires a virtual setup: here we use Pardon’s construction of Hamiltonian-Floer cohomology for a general symplectic manifold via his VFC theory [Par16], as symplectic cohomology is just a special case of Hamiltonian-Floer cohomology. The latter concern requires understanding the behavior of contact dynamics and holomorphic curves under connected sums. On the other hand, to infer information on the topology of the filling, it is sufficient to use the tautological exact sequence and the restriction on the connecting homomorphism for filtered symplectic cohomology (which involves Reeb orbits up to an action threshold), despite the fact that this is not an invariant of the contact manifold (as it depends on the contact form):
In particular, we make essential use of specific models for the Reeb dynamics, and one can think of this as arguing with cohomology on the chain level rather than with the invariant directly. In particular, our arguments involve careful computations of various boundary maps on the nose.
4.1. Main result
@@
Theorem 4.1 (Homological Constraints on Fillings).
Let $W_{0}$ be a $(2n+2)$-dimensional flexible Weinstein domain with $n\geq 2,$ and let $(M,\xi)$ be any contact manifold.
  1. (a)
    Suppose that $c_{1}(W_{0})$ and $c_{1}(\xi)$ both vanish. If $W$ is a strong filling of $Y=M\#\partial W_{0}$, then the kernel of $H_{n}(\partial W_{0};\mathbb{Q})\to H_{n}(W_{0};\mathbb{Q})$ contains the kernel of
    $H_{n}(\partial W_{0};\mathbb{Q})\stackrel{{\scriptstyle 0\oplus\operatorname{Id}}}{{\hookrightarrow}}H_{n}(M;\mathbb{Q})\oplus H_{n}(\partial W_{0};\mathbb{Q})=H_{n}(M\#\partial W_{0};\mathbb{Q})\rightarrow H_{n}(W;\mathbb{Q}).$
  2. (b)
    Let $W_{0}$ be subcritical (without any assumption on $c_{1}=0$). Then, for any strong filling $W$ of $Y$ we have that $H_{n}(\partial W_{0};\mathbb{Q})\to H_{n}(W;\mathbb{Q})$ is injective.
Note that, in the case that $W_{0}$ is subcritical, the first claim reduces to the fact that $H_{n}(\partial W_{0};\mathbb{Q})\to H_{n}(W;\mathbb{Q})$ is injective. Indeed, the first map $H_{n}(\partial W_{0};\mathbb{Q})\to H_{n}(W_{0};\mathbb{Q})$ is automatically injective in this case: this is a consequence of the long exact sequence of the homology of the pair, as $H_{n+1}(W_{0},\partial W_{0};\mathbb{Q})$ is generated by the cocores of the Weinstein handles of index $n+1$, and hence there are none if $W_{0}$ is subcritical. The non-trivial content of statement (b) is that subcriticality implies that $H_{n}(\partial W_{0};\mathbb{Q})\to H_{n}(W;\mathbb{Q})$ is injective without the assumption that $c_{1}(\xi)$ and $c_{1}(W_{0})$ vanish.
@@
Theorem 4.1 is stated in this way only for the purpose of constructing exotic contact structures on spheres in this paper, but in fact, a stronger version of this result, giving homological information on the fillings in some other degrees for $W_{0}$ flexible, can also be obtained by similar arguments. We also point out that, when $M=\emptyset$ and $W_{0}$ is subcritical, one can use a strategy of filling by pseudo-holomorphic curves similar to [BGZ19] to obtain a similar result, which was used by the first three authors [BGM22, Theorem B] to obstruct strong fillings.
In the proof of Theorem 4.1, it is important to rule out the effects from Reeb orbits on $M$ as well as the boundary of the Weinstein $1$-handle in the connected sum. The former can be easily dealt with if we scale up the contact form on $M$, and those Reeb orbits will have large periods and hence not be seen, as we will use a fixed period threshold for filtered symplectic cohomology on $\partial W_{0}$. It is the Reeb orbits on the handle that divides Theorem 4.1 into two cases. In the general flexible case, the Reeb orbits created on the handle will have sufficiently high Conley-Zehnder indices, see e.g. Lemma 4.3. Those orbits do not affect the piece of symplectic cohomology we care about. This uses the $\mathbb{Z}$-grading in an essential way. In the subcritical case, we make use of the splitting form and a special choice of almost complex structure so that we can rule out the interaction with those orbits on the handle by an energy argument, which is inspired by similar considerations in the work of Chiang-Ding-van Koert [CDvK16].